# Busting_the_princelings_replication.bib Are Princelings Truly Busted? Evaluating Transaction Discounts in China’s Land Market

### v1

@@ 3 Misspecifying Area
For their later tables, CK filter and transform the land transaction records to group by relevant features. When they aggregate transaction records at the firm and year level—such that they list the total area of all of firm ’s transactions in a given year—the value that they define as the natural logarithm of the area purchased is actually the area purchased divided by 1,000,000. However, CK interpret these area/1,000,000 values as logarithms; I include specific examples in the Appendix. Thus, I re-interpret their coefficients for the tables impacted by this transformation error, for cases with and without duplicates. This error extends to CK’s Tables VI, VIII, IX, XI, and XII as well as Figure VI.101010The tables discussed previously (CK’s Tables V and X) do not suffer from these issues because they do not contain the aggregated measures of the logarithm of area, remaining at the parcel-level. The remaining tables (I, II, III, IV, and VII) offer summary information and are not impacted by the misinterpretation; further, of CK’s figures, only Figure VI is impacted by the interpretation error.
@@ 3.1 Replicating CK’s Table VI
CK’s Table VI “Quantity of Land Purchased by Princeling Firms, 2004-2016” regresses the area of land purchased in the primary land market, divided by 1,000,000, on the dummy indicator signaling a princeling connection, replicated in Panel A of Table 4.111111While CK state that industry fixed effects are included in Table VI, they are omitted from the calculation. While CK originally find that princeling-connected firms purchased 0.2% more land each year than their nonprinceling counterparts, the regression coefficient actually suggests that princeling-connected firms purchased an additional 2,000 square meters of land annually compared to their nonprinceling peers, across all land transactions combined. Contextually, 90.8% of the rows have a total annual area purchased of zero square meters, as few firms (only 14%) purchase land in multiple years between 2004 and 2016. Excluding the rows in the panel that have zero area, the median area of land purchased annually by a firm is 26,668 square meters; the mean is 68,193 square meters, as a small number of firms have very high yearly transaction amounts (up to 14.67 million square meters). I thus find that this median more accurately reflects the average firm’s yearly transaction volume. When the rows with zero area are included, the median is 0 square meters, and the mean is 6,250 square meters. Excluding zeros, the average number of parcels purchased annually per firm is 1, although the maximum is 826 parcel purchases.
@@
The 2,000 square meter advantage is therefore a 7.5% increase over the median parcel size (excluding rows with zero area), much more than CK’s original 0.2%. In column (2), firms connected to Politburo Standing Committee members purchase 32,000 square meters of additional land per year compared to nonprincelings—more than doubling the median parcel size; this is far above the 3% that CK originally obtained (column (2), exp(0.001+0.031)). Column (3) indicates that retirement has a negligible effect on the quantity of land purchased.
@@
When Table VI is recalculated without duplicates, as highlighted in Panel B of Table 4, the result of column (1) does not change, but column (2) now suggests that princelings with Politburo Standing Committee connections purchase 41,000 additional square meters of land a year than their nonprinceling peers. The retirement of an official (column (3)) is now associated with a significant increase in the area of land purchased by princelings, with princeling firms connected to retired officials purchasing, on average, an additional 59,000 square meters of land—bringing their total purchases to more than three times the median parcel size.121212Note that a relatively small number of firms are connected to retired officials (668), and the data suggests that among purchases made by firms with connections to retired officials, smaller parcels are disproportionately represented among the duplicates, causing this large increase in the coefficient. Of the unique transactions involving retired officials, 46% have at least one duplicate, with some having as many as 22 duplicates. Indeed, the mean and median yearly parcel purchase areas are significantly higher for firms connected to retired officials while those for nonprincelings and for princelings without this connection are roughly equal to each other.
| Area of land purchased/1,000,000 | ||||||
@@ 3.2 Tables VIII, IX, and XII
For Tables VIII, IX, and XII, I am unable to replicate CK’s aggregated area measures. These specifications focus on the effect of business favor on promotion, examining whether there is an association between officials selling land to princelings for discounted prices and their likelihood of promotion. As such, they regress political turnover for provincial officials on different measures of princeling connections (Table VIII), political turnover for municipal officials (Table IX), and then political turnover for both types of officials (Table XII); they thus aggregate the total land sold to firms under each official’s purview by year and province/municipality, respectively. CK likely have another variable that connects a transaction to an official at the province/municipality level, but they exclude this linking variable from the replication data, preventing recalculation. Yet, in all cases, the logarithm of the area of land sold seems extremely small, suggesting that the “logarithm of area sold” is again area/1,000,000.131313For Table XII, CK utilize the logarithm of area as the independent variable for a set of regressions, calling it , yet based on the description of this variable as measuring the “quantity of land purchased,” it is not clear whether they mean to interpret it as the logarithm of area or as a scaled measure of area (p. 220). Regardless, as CK do not offer a detailed interpretation of these coefficients, the substantive implications do not change: the effect of on the promotion of party secretaries drops for transactions occurring after 2012 for provincial and municipal secretaries and in regions with central inspections for provincial party secretaries.
@@ 3.3 Table XI and Figure VI
CK’s Table XI investigates whether the quantity of land purchased by princeling firms changed after President Xi took office and began the anticorruption campaign.141414Note that one of Xi’s first initiatives when taking over as party leader was the anticorruption campaign—he announced at the 18th National Congress (the top meeting for setting agendas within the party) that he would work to catch the “tigers” (high-ranking corrupt officials) and “flies” (low-ranking corrupt officials) to ensure the party’s survival ([wong_new_2012]; [xi_power_2013]; [yuen_disciplining_2014]). Table XI thus regresses the land area purchased/1,000,000 on whether a transaction occurred after 2012 (column (1)), whether a central inspection took place (column (2)), and whether the transaction occurs in a province wherein Xi replaced the party secretary; these results are illustrated in Table 5. CK originally found that princeling firms purchased more land than their nonprinceling counterparts by approximately 8.1%, but this advantage dropped by 2.3% after 2012. Reinterpreting these coefficients yields that princeling firms annually purchased 78,000 more square meters of land than their nonprinceling counterparts, on average, before 2012, dropping to 56,000 square meters afterwards. With the median total purchased by a firm in a given year and province being 26,114 square meters, these purchase advantages reflect a sizeable princeling benefit, with the princeling advantage suggesting that princelings purchased almost quadruple the median land purchase volume before 2012 and then only triple the median after 2012.151515The panel data used in Table XI lists each firm’s parcel purchases grouped by each province and year between 1991-2016. Given few firms make frequent land purchases, 95.4% of the observations have zero square meters of land purchased in a given province/year. The reported median includes only those nonzero transactions as the median would be 0 otherwise. The mean is 3,076 square meters when including rows with zero transactions per year/province and 66,423 square meters when these zeros are excluded.
| Area of land purchased/1,000,000 | ||||||||
@@
In column (4), CK originally find that the binary variable reflecting whether a transaction occurs after 2012 becomes insignificant when the dummy variables for central inspection and Xi-appointed officials are included. They further note that “given that the sum of the coefficients in column (4) is not significantly different from zero, these two measures together effectively eliminate the advantage of the princeling firms in purchasing a larger quantity of land” (p. 219).161616Here, CK are summing the coefficients that include —that is, the dummy variable and its interaction terms. This result remains in the absence of the intended logarithm.
When duplicates are removed (Table 5, Panel B), the magnitude of the princeling effect drops across all specifications but remains statistically significant: column (1), for instance, indicates that princeling firms purchased 48,000 square meters more land than their nonprinceling counterparts annually. After 2012, this effect does not change in a statistically significant way. For column (4), as above, “transactions after 2012” is insignificant when “central inspection” and “Xi-appointed officials” are included, the coefficient changing sign now that duplicates have been removed, and the sum of the princeling dummy variable and its interacted terms is not statistically different from zero. “Transactions after 2012” is thus no longer significant in any specification, suggesting that the effect of the anticorruption campaign may not be as clear-cut.
@@ 4 Discussion
The presence of duplicates and the misinterpreted logarithm of area raise important implications about CK’s original results, chiefly the size of the princeling effect in terms of square meters. While the princeling effect often remained statistically significant but shrunk when duplicates were removed (as apparent in Table 5), the economic significance is not fully clear: given the significant variation in firm parcel purchase volumes and sizes, a benefit of around 78,000 square meters (Table 5, column (1)) is extremely significant for a firm whose yearly purchase total per province is around the median of 26,114 square meters, but less so for one whose yearly purchase volume is over 4 million square meters (as is the case of the princeling with the largest land purchase total in a single year).
Importantly, these results also raise questions about underlying data quality; with the panel data used in CK’s Tables VI and XI having approximately 91% and 95% zeros, respectively, in the area column, effects are estimated from a small number of nonzero coefficients and are sensitive to duplicates. Given these consistency issues, taking the logarithm of the area would certainly afford more meaningful insight. Yet, when the actual logarithm of area is used (as described in the Appendix), the results are unstable, changing sign and direction across specifications.
Further, the causal implications of the anticorruption campaign are limited when duplicates are removed, with Figure 1(b) highlighting that the difference in princeling and nonprinceling land purchase volumes dropped before 2012. Similarly, Table 5 confirms that the coefficient for a transaction occurring after 2012 changes sign and is no longer statistically significant—although the variables representing central inspections and Xi-appointed officials remain significant—perhaps suggesting that the anticorruption campaign depends more on regional enforcement and signaling. Given these results as well as the instability of the analysis when the true logarithm is used, further investigation should be done to determine whether the anticorruption campaign truly plays a causal role in reducing princeling discounts.
5 Appendix A: Duplicate examples
@@
6 Appendix B: Misspecifying area/1,000,000 as log(area)
@@
In the year 2015, Firm 1 has only one transaction with a raw area of 17,795.3 square meters. Thus, when CK aggregate the total amount of area sold to each firm in each year, Firm 1’s total for 2015 is still just 17,795.3 square meters. Yet, in their dataset, the value that CK have for the reported is 0.01764, which is approximately 17,795.3/1,000,000. The slight variation in numbers seems to come from their calculation: they likely take the logarithm of the original area divided by 100. They then truncate/round it, exponentiate it, and divide it by 10,000.
@@
A second example with aggregation confirms that this is the case. For instance, firm 83 has 2 land transactions that occur in 2013. I include both of these transaction records, as well as the aggregated entry, in the two panels of Table 7.
@@
In these cases and all others, CK’s reported is approximately area/1,000,000, the slight difference again resulting from CK’s rounding during intermediate steps. In Section 3, I reinterpret CK’s main tables in light of the fact that area is scaled rather than logged; the tables recalculated with correct log values are included in Appendix C.
7 Appendix C: Calculations utilizing corrected log(area) values
In this Appendix, I conduct Chen and Kung’s calculations with the correct logarithm of area purchased and find, in many cases, substantially different results, with and without duplicates. In particular, the princeling effect changes its sign and direction, differing in cases with and without duplicates; in multiple instances, the results are qualitatively different when the correct logarithm of area value is used, even when calculated on the original data. Further, there is evidence of a sharp drop in the princeling advantage prior to the anticorruption campaign. As above, I regenerate Tables VI and XI as well as Figure VI, which are impacted by these errors. For this analysis, CK seek to analyze the differences in quantity purchased between princelings and nonprincelings overall, not just those firms who buy land in a given year—and presumably, princelings may be more likely to purchase land in a given year in the first place. Thus, I calculate log(area), not using a scaling factor, and following CK, I add 1 to the area value to “deal with observations of 0” prior to taking the log (p. 212).
For Table VI, CK originally interpret the regression coefficient as suggesting that princeling-connected firms purchased 0.2% (column (1), exp(0.002)) more land each year than their nonprinceling counterparts; they find the quantity of land purchased by firms connected to Politburo Standing Committee members (PSCM) is 3% higher (column (2), exp(0.001+0.031)) than that purchased by nonprincelings and that retirement has a negligible effect on the quantity of land purchased by princeling firms (column (3)). Yet, as highlighted in Panels A and B of Table 8, these results differ significantly from calculations that involve the correct logarithm of area values. In the baseline specification, princeling-connected firms purchase 12.4% more land each year than their nonprinceling counterparts. The other specifications are more complicated: column (2) suggests that princeling firms purchase 26.2% less land than their nonprinceling counterparts but that princeling firms with PSCM connections purchase a massive 179,084% more land (exp(0.304+7.795)1) than nonprinceling firms. Similarly, column (3) indicates that princelings purchase 6.1% less land than their nonprinceling counterparts but that princeling firms connected to retired officials purchase 344,511% more land (exp(0.063+8.208)1) than their nonprinceling
@@ 7 Appendix C: Calculations utilizing corrected log(area) values
| Princeling firms*PSCM | 7.795*** | 7.208*** | ||||
@@
| Princeling firms*Retired | 8.208*** | 7.825*** | ||||
@@
peers. Such percentages are so large as to be nonsensical: with the median area of land purchased annually by a firm being 26,668 square meters, a 344,511% increase above this would be 91,900,739 square meters while a 179,084% increase would be 47,784,870 square meters—and these are well above the maximum annual firm purchase volume of around 14.7 million square meters.171717As above, this median statistic includes only the rows in the panel that have nonzero area values (pre-log); otherwise, the median would be 0. The mean is 6,250 square meters with zeros and 68,193 square meters excluding zeros, as a small number of firms have very high transaction amounts and skew the mean upwards. I thus find the median more accurately reflects the average firm’s yearly transaction volume but note that the princeling advantage would yield even more significant increases in area if the mean were used instead. These specifications are also interesting because columns (2) and (3) suggest that regular princelings purchase less land than nonprincelings, but princelings with PSCM and retired official connections purchase significantly more, reversing CK’s original conclusions—and the regressions calculated without duplicates (Panel B, columns (1)-(3)) further echo these results.
CK’s Table XI investigates whether the quantity of land purchased by princeling firms changed after President Xi took office and began the anticorruption campaign. Table XI thus regresses the logarithm of land area purchased on whether a transaction occurred after 2012 (column (1)), whether a central inspection took place (column (2)), and whether the transaction occurred in a province wherein Xi replaced the party secretary (column (3)). CK originally found that princeling firms purchased more land than their nonprinceling counterparts by approximately 8.1%; after 2012, princeling land purchases dropped by more than 2.3% (and by 29% in their purchase quantity premium compared to nonprincelings). With the corrected log area, as shown in Panel A of Table 9, the regressions suggest princeling firms purchased 244,694% (exp(7.803)1) more land than their nonprinceling counterparts prior to 2012, and after 2012, their land purchases increase by 52,112% (exp(7.803)exp(7.803+0.193)), or a 21.3% increase (() in their purchase quantity premium compared to nonprincelings.
@@
For the interacted central inspection equation (column (2)), the estimation suggests that the land sold to princeling firms increased by 203,501% (exp(7.938)exp(7.938+9.0148.468)), or that their purchase quantity premium increased by 72.7% compared to nonprinceling firms (). These percentages are so large as to be wholly nonsensical: with the median total purchased by a firm in a given year and province being 26,114 square meters, these numbers suggest the average princeling advantage would be 63,925,369 square meters (per column (1)’s 244,694% growth)—a value far above the maximum annual land purchase by a firm.181818As noted in the main text, many firms purchase land infrequently, and 95.4% of the observations have zero square meters purchased in a given province/year. The reported median includes only those nonzero transactions (as the median would otherwise be 0). The mean is 3,076 square meters when including rows with zero square meters sold and 66,423 square meters when excluding zeros. Likewise, in CK’s original analysis, they find that the binary variable reflecting whether a transaction occurs after 2012 becomes insignificant in column (4) when the dummy variables for central inspection and Xi-appointed officials are included, but when recalculated, this is no longer the case.
@@
When duplicate rows are removed, as highlighted in Panel B of Table 9, the coefficients are still quite large: Panel B’s column (1), for instance, indicates that princeling firms purchased more land than their nonprinceling counterparts by 137,371% (exp(7.226)1). After 2012, their land purchases increase by 128,507% (exp(7.226)exp(7.226+0.660)), which is a 93.5% increase () in their purchase quantity premium relative to nonprincelings. Yet, due to the extremely large percentages of princeling advantage in purchase quantity premium, the results are difficult to interpret and nonsensical, likely reflecting high variability in parcel sizes, the need for further controls, the relatively small number of princeling firms in the overall data, and the sensitivity resulting from adding 1 to area prior to logging it. These results are also interesting because they run counter to that from the recalculated Table VI, which suggested princeling firms—in the presence of interacted terms—purchased less land than their nonprinceling counterparts, confirming these data and estimation issues.

### v2

@@ 3 Misspecifying Area
For their later tables, CK filter and transform the land transaction records to group by relevant features. When they aggregate transaction records at the firm and year level—such that they list the total area of all of firm ’s transactions in a given year—the value that they define as the natural logarithm of the area purchased in their paper is actually the area purchased () divided by 1,000,000, per their code. However, CK interpret these area/1,000,000 values as logarithms; I include specific examples in the appendix. Thus, I re-interpret their coefficients for the tables impacted by this transformation error, for cases with and without duplicates. This error extends to CK’s Tables VI, VIII, IX, XI, and XII as well as Figure VI.999The previously discussed Tables V and X do not suffer from these issues because they do not contain aggregated measures of the logarithm of area, remaining at the parcel-level. The remaining tables (I, II, III, IV, and VII) offer summary information and are not impacted by the misinterpretation; of CK’s figures, only Figure VI is impacted.
@@ 3.1 Replicating CK’s Table VI
CK’s Table VI “Quantity of Land Purchased by Princeling Firms, 2004-2016” regresses the area of land purchased in the primary land market, divided by 1,000,000, on the dummy indicator signaling a princeling connection, replicated in Panel A of Table 4.101010While CK state that industry fixed effects are included in Table VI, they are omitted from their code. While CK originally find that princeling-connected firms purchased 0.2% ((exp(0.002) more land each year than their nonprinceling counterparts, the regression coefficient actually suggests that princeling-connected firms purchased an additional 2,000 (1,000,000) square meters of land annually compared to their nonprinceling peers, across all land transactions combined. Contextually, 90.8% of rows have a total annual area purchased of zero square meters, as few firms (only 14%) purchase land in multiple years between 2004 and 2016. Excluding firms who purchase zero area, the median area of land purchased annually by a firm is 26,668 square meters; the mean is 68,193 square meters, as a small number of firms have very high yearly transaction amounts (up to 14.67 million square meters). I thus find that the median more accurately reflects the average firm’s yearly transaction volume. When the rows with zero area are included, the median is 0 square meters, and the mean is 6,250 square meters. Excluding zeros, the average number of parcels purchased annually per firm is 1, although the maximum is 826 parcel purchases.
@@
The 2,000 square meter advantage is therefore a 7.5% (2,000/26,668) increase over the median parcel size (excluding rows with zero area), much more than CK’s original 0.2%. In column (2), firms with PSCM connections purchase 32,000 square meters ((0.001+0.031) of additional land per year compared to nonprincelings—more than doubling the median parcel size; this is far above the 3% that CK originally obtained (column (2), (exp(0.001+0.031)). Column (3) indicates that retirement has a negligible effect on the quantity of land purchased.
@@
When Table VI is recalculated without duplicates (Table 4, Panel B), the result of column (1) does not change, but column (2) now suggests that princelings with PSCM connections purchase 41,000 (0.0411,000,000) additional square meters of land a year than their nonprinceling peers. Retired official connections (column (3)) are now associated with a significant increase in the area of land purchased by princelings, with princeling firms connected to retired officials purchasing, on average, an additional 59,000 ((0.001+0.058)1,000,000) square meters of land relative to nonprincelings—bringing their total purchases to more than three times the median parcel size.111111Note that a relatively small number of firms (668) are connected to retired officials, and the data suggests that among purchases made by firms with connections to retired officials, smaller parcels are disproportionately represented among the duplicates, causing this large increase in the coefficient. Of the unique transactions involving retired officials, 46% have at least one duplicate, with some having as many as 22 duplicates. Indeed, the mean and median yearly parcel purchase areas are significantly higher for firms connected to retired officials while those for nonprincelings and for princelings without this connection are roughly equal to each other.
| Area of land purchased ()/1,000,000 | ||||||
@@ 3.2 Tables VIII, IX, and XII
For Tables VIII, IX, and XII, I am unable to replicate CK’s aggregated area measures. These specifications focus on whether there is an association between officials selling land to princelings for discounted prices and their likelihood of promotion. As such, they regress political turnover for provincial officials on different measures of princeling connections (Table VIII), political turnover for municipal officials (Table IX), and then political turnover for both types of officials (Table XII); they thus aggregate the total land sold to firms under each official’s purview by year and province/municipality, respectively. CK likely has another variable that connects a transaction to an official at the province/municipality level, but they exclude this linking variable from the replication data, preventing recalculation.121212[5] was able to replicate part of CK’s Table IX, specifically the coefficient on GDP growth in columns (7) and (8); Wiebe noted that he received additional data from one of the authors. Yet, in all cases, the logarithm of the area sold seems extremely small, suggesting that this “logarithm of area sold” is again area () /1,000,000.131313For Table XII, CK utilize the logarithm of area as the independent variable for a set of regressions, calling it , yet as this variable is described as measuring the “quantity of land purchased,” it is not clear whether they mean to interpret it as the logarithm of area or as area in square meters (p. 220). Regardless, as CK do not offer a detailed interpretation of these coefficients, the substantive implications do not change: the effect of on the promotion of party secretaries drops for transactions occurring after 2012 for provincial and municipal secretaries and in regions with central inspections for provincial party secretaries.
@@ 3.3 Table XI and Figure VI
CK’s Table XI investigates whether the quantity of land purchased by princeling firms changed after President Xi took office and began the anticorruption campaign.141414One of Xi’s first initiatives as party leader was the anticorruption campaign: he announced at the 18th National Congress (the top meeting for setting agendas within the party) that he would work to catch the “tigers” (high-ranking corrupt officials) and “flies” (low-ranking corrupt officials) to ensure the party’s survival ([6]; [7]; [8]). Table XI thus regresses the land area purchased () /1,000,000 on whether a transaction occurred after 2012 (column (1)), whether a central inspection took place (column (2)), and whether the transaction occurs in a province wherein Xi replaced the party secretary (column (3)); these results are illustrated in Table 5. CK originally found that princeling firms purchased more land than their nonprinceling counterparts by approximately 8.1% ((exp(0.078)), but this advantage dropped by 2.3% ((exp(exp())) after 2012. Reinterpreting these coefficients yields that princeling firms annually purchased 78,000 (1,000,000) more square meters of land than their nonprinceling counterparts, on average, before 2012, dropping to 56,000 (1,000,000) square meters afterwards. With the median total firm purchase in a year/province being 26,114 square meters, these purchase advantages reflect a sizeable princeling benefit, suggesting that princelings purchased almost quadruple the median land purchase volume before 2012 and then only triple the median after 2012.151515The panel data used in Table XI lists each firm’s parcel purchases grouped by each province and year between 1991-2016. Given few firms make frequent land purchases, 95.4% of the observations have zero square meters of land purchased in a given province/year. The reported median includes only those nonzero transactions as the median would be 0 otherwise. The mean is 3,076 square meters when including rows with zero transactions per year/province and 66,423 square meters when these zeros are excluded.
In column (4), CK originally find that the binary variable reflecting whether a transaction occurs after 2012 becomes insignificant when the dummy variables for central inspection and Xi-appointed officials are included. They further note that “given that the sum of the coefficients in column (4) is not significantly different from zero, these two measures together effectively eliminate the advantage of the princeling firms in purchasing a larger quantity of land” (p. 219).161616Here, CK sum all coefficients including —that is, the dummy variable and its interactions. This result remains in the absence of the intended logarithm.
When duplicates are removed (Table 5, Panel B), the magnitude of the princeling effect drops across all specifications but remains statistically significant: column (1), for instance, indicates that princeling firms purchased 48,000 (1,000,000) square meters more land than nonprincelings annually. After 2012, this effect does not change in a statistically significant way. For column (4), as above, “transactions after 2012” is insignificant when “central inspection” and “Xi-appointed officials” are included, the coefficient changing sign now that duplicates have been removed, and the sum of the princeling dummy variable and its interacted terms is not statistically different from zero. “Transactions after 2012” is thus no longer significant in any specification, suggesting that the effect of the anticorruption campaign may not be as clear-cut.
| Area of land purchased ()/1,000,000 | ||||||||
@@ 4 Discussion
The presence of duplicates and the misinterpreted logarithm of area raise important implications about CK’s original results: while the princeling effect often remained statistically significant but shrunk when duplicates were removed (as apparent in Table 5), the economic significance is not fully clear. Given the significant variation in firm parcel purchase volumes and sizes, a benefit of around 78,000 square meters (Table 5, column (1)) is extremely significant for a firm whose yearly purchase total per province is around the median of 26,114 square meters, but less so for one whose yearly purchase volume is over 4 million square meters (the year-level maximum).
Importantly, these results also raise questions about underlying data quality; with the panel data used in CK’s Tables VI and XI having approximately 91% and 95% zeros, respectively, in the area column, effects are estimated from a small number of nonzero coefficients and are sensitive to duplicates. Given these consistency issues, taking the logarithm of the area would certainly afford more meaningful insight. Yet, when the logarithm of area is used (as described in the appendix), the results are unstable, changing sign and direction across specifications.171717In the appendix, the logarithm is calculated as log(y+1), as this is the method that CK utilize in their paper. Given log(y+1) assigns an arbitrary weight to the extensive margin effect, I also estimate these specifications using Poisson regression and still find unstable effect estimates. I further examine the extensive margin effect via logistic regression and a linear probability model.
Further, the causal implications of the anticorruption campaign are limited when duplicates are removed, with Figure 1(b) highlighting that the difference in princeling and nonprinceling land purchase volumes dropped before 2012. Similarly, Table 5 confirms that the coefficient for a transaction occurring after 2012 changes sign and is no longer statistically significant—although the variables representing central inspections and Xi-appointed officials remain significant—perhaps suggesting that the anticorruption campaign depends more on regional enforcement and signaling. Given these results as well as the instability of the analysis when the logarithm is used (per the appendix), further investigation should be done to determine whether the anticorruption campaign truly plays a causal role in reducing princeling discounts.
Appendix
A.1 Duplicate examples
@@
A.2 Misspecifying area/1,000,000 as log(area)
@@
In the year 2015, firm 1 has only one transaction with a raw area of 17,795.3 square meters. Thus, when CK aggregate the total amount of area sold to each firm in each year, firm 1’s total for 2015 is still just 17,795.3 square meters. Yet, in their dataset, the value that CK has for the reported is 0.01764, which is approximately 17,795.3/1,000,000. The slight variation in numbers seems to come from their calculation: they likely take the logarithm of the original area divided by 100. They then truncate/round it, exponentiate it, and divide it by 10,000.
@@
A second example with aggregation confirms that this is the case. For instance, firm 83 has 2 land transactions that occur in 2013. I include both of these transaction records, as well as the aggregated entry, in the two panels of Table A2.
@@
In these cases and all others, CK’s reported is approximately area/1,000,000, the slight difference again resulting from CK’s rounding during intermediate steps. In Section 3, I reinterpret CK’s main tables in light of the fact that area is scaled rather than logged; the tables recalculated with correct log values are included in Appendix A.3.
A.3 Estimating semi-elasticities
In this section, I estimate what Chen and Kung claim to have done in their paper—that is, conduct the calculations with the logarithm of area purchased—and find, in many cases, substantially different results, with and without duplicates. In particular, the princeling effect changes its sign and direction, differing in cases with and without duplicates; in multiple instances, the results are qualitatively different when the logarithm of area value is used, even when calculated on the original data. Further, there is evidence of a sharp drop in the princeling advantage prior to the anticorruption campaign. As above, I regenerate CK’s Tables VI and XI as well as Figure VI, which are impacted by these errors. For this analysis, CK seek to analyze the differences in quantity purchased between princelings and nonprincelings overall, not just those firms who buy land in a given year—and presumably, princelings may be more likely to purchase land in a given year in the first place. Thus, I follow CK in calculating log(area + 1), not using a scaling factor and adding 1 to the area value to “deal with observations of 0” prior to taking the log (p. 212). However, because this approach assigns an arbitrary weight to the extensive margin effect, I also recalculate these tables using Poisson regression in Section A.3.1. I then evaluate the extensive margin effect in Section A.3.2.
For Table VI, CK originally interpret the regression coefficient as suggesting that princeling-connected firms purchased 0.2% (column (1), (exp(0.002) more land each year than
@@ A.3 Estimating semi-elasticities
| Princeling firms*PSCM | 7.795*** | 7.241*** | ||||
@@
| Princeling firms*Retired | 8.208*** | 7.834*** | ||||
@@
their nonprinceling counterparts; they find that the quantity of land purchased by firms connected to Politburo Standing Committee members (PSCM) is 3% higher (column (2), (exp(0.001+0.031)) than that purchased by nonprincelings and that retirement has a negligible effect on the quantity of land purchased by princeling firms (column (3)). Yet, as highlighted in Panels A and B of Table A3, these results differ significantly from calculations that involve the actual logarithm of area values. In the baseline specification, princeling-connected firms purchase 12.4% ((exp() more land each year than their nonprinceling counterparts. The other specifications are more complicated: column (2) suggests that princeling firms purchase 26.2% less land ((exp() than their nonprinceling counterparts but that princeling firms with PSCM connections purchase a massive 179,084% more land ((exp(0.304+7.795)1)) than nonprinceling firms. Similarly, column (3) indicates that princelings purchase 6.1% less land ((exp() than their nonprinceling counterparts but that princeling firms connected to retired officials purchase 344,511% more land ((exp(0.063+8.208)1) than their nonprinceling peers. Such percentages are so large as to be nonsensical: with the median area of land purchased annually by a firm being 26,668 square meters, a 344,511% increase above this would be 91,900,739 square meters (26,668exp()) while a 179,084% increase would be 47,784,870 square meters (26,668exp(0.304+7.795))—and these are well above the maximum annual firm purchase volume of around 14.7 million square meters.181818As above, this median statistic includes only the rows in the panel that have nonzero area values (pre-log); otherwise, the median would be 0. The mean is 6,250 square meters with zeros and 68,193 square meters excluding zeros, as a small number of firms have very high transaction amounts and skew the mean upwards. I thus find the median more accurately reflects the average firm’s yearly transaction volume but note that the princeling advantage would yield even more significant increases in area if the mean were used instead. These specifications are also interesting because columns (2) and (3) suggest that regular princelings purchase less land than nonprincelings, but princelings with PSCM and retired official connections purchase significantly more, reversing CK’s original conclusions—and the regressions calculated without duplicates (Panel B, columns (1)-(3)) further echo these results.
CK’s Table XI investigates whether the quantity of land purchased by princeling firms changed after President Xi took office and began the anticorruption campaign. Table XI thus regresses the logarithm of land area purchased on whether a transaction occurred after 2012 (column (1)), whether a central inspection took place (column (2)), and whether the transaction occurred in a province wherein Xi replaced the party secretary (column (3)). CK originally found that princeling firms purchased more land than their nonprinceling counterparts by approximately 8.1%; after 2012, princeling land purchases dropped by more than 2.3%. When using log(area+1), as shown in Panel A of Table A4, the regressions suggest princeling firms purchased 244,694% ((exp(7.803)1)) more land than their nonprinceling counterparts prior to 2012, and after 2012, their land purchases increase by 52,112 percentage points ((exp(7.803+0.193)exp(7.803))).
@@
For the interacted central inspection equation (column (2)), the estimation suggests that the land sold to princeling firms increased by 203,501 percentage points ((exp(7.938+9.0148.468)exp(7.938)). These percentages are so large as to be wholly nonsensical: with the median total purchased by a firm in a given year and province being 26,114 square meters, these numbers suggest the average princeling advantage would be 63,925,369 square meters (per column (1)’s 244,694% growth, as 26,114 exp(7.803—a value far above the maximum annual land purchase by a firm.191919As noted in the main text, many firms purchase land infrequently, and 95.4% of the observations have zero square meters purchased in a given province/year. The reported median includes only those nonzero transactions (as the median would otherwise be 0). The mean is 3,076 square meters when including rows with zero square meters sold and 66,423 square meters when excluding zeros. Likewise, CK’s original analysis finds that the binary variable reflecting whether a transaction occurs after 2012 becomes insignificant in column (4) when the dummy variables for central inspection and Xi-appointed officials are included, but when recalculated, this is no longer the case.
@@
When duplicate rows are removed, as highlighted in Panel B of Table A4, the coefficients are still quite large: Panel B’s column (1), for instance, indicates that princeling firms purchased more land than their nonprinceling counterparts by 139,449% ((exp(7.241)1)). After 2012, their land purchases increase by 127,763 percentage points ((exp(7.241+0.650)exp(7.241))). Yet, due to the extremely large percentages of princeling advantage, the results are difficult to interpret and nonsensical, likely reflecting high variability in parcel sizes, the need for further controls, the relatively small number of princeling firms in the overall data, and the sensitivity resulting from adding 1 to area prior to logging it. These results are also interesting because they run counter to that from the recalculated version of CK’s Table VI (Table A3), which suggested that princeling firms—in the presence of interacted terms—purchased less land than their nonprinceling counterparts, confirming these data and estimation issues.